Find the absolute maximum and minimum values of f(x,y) = 2x + y4 on the set D = {(x,y) x2 + y2 <1}.
Answer #1

f(x, y) = 2x + y4 f = 0 6x²=0x=0 f =0 4 y = 0 => y=0 (0:0) is critical point it is inside circle now on boundrary.x2 + y2 =1

1 9 = 2x=+= 8 16 II 1 9 + 4 16 2 13 = 16 13 f = 2 x -+ 2 2 8 9 16 13 16 x=0 then y=+1 y= 0x==1 f(1,0)=2,5(-1,0)=-2 f(0,1)=1 f